A gas expands and does 82.7 J of work on the surroundings. At the same time, the gas loses 103 J of energy to the surroundings as heat. What is the change in the internal energy of the gas?
Would this be ΔE = q + w (-103) + (82.7) = -20.3. I get confused with the positive and negative charges.
1
Expert's answer
2015-10-19T02:55:35-0400
Internal energy of gas ΔU = Q + W consists of work that is done and heat that is lost.
Gas loses 82.7 J of energy in form of work and 103 J in form of heat.
Finally,
ΔU = Q + W = -82.7 – 103 = -185.7 J
Gas loses 185.7 J of energy.